OpenStudy (anonymous):

Arrange the list of numbers in increasing order. 2/5, 4/9, 3/7

OpenStudy (anonymous):

2/5 3/7 4/9

OpenStudy (anonymous):

divide the numbers and see the value of the answers

OpenStudy (anonymous):

2/5= .4 3/7= .42 4/9= .44

OpenStudy (anonymous):

Get the Denominator Joint . It can be 315 :) Then we have : 2/5 = x/315 4/9 = y/315 3/7=z/315

hero (hero):

2/5 and 4/9 Cross multiply (2)(9) < (5)(4) so 2/5, and 4/9 are already in the right order Now cross multiply 4/9 and 3/7 (4)(7) > (9)(3) Since 28 > 27 swap the order of 4/9 and 3/7 to get: 2/5, 3/7, 4/9 Now test 2/5 and 3/7 (2)(7) < (5)(3) Therefore 2/5, 3/7, 4/9 is the correct order

OpenStudy (anonymous):

aka 126/315, 135/315, 140/315

hero (hero):

It's easier to just cross multiply and swap

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

2/5 =0.40 4/9 = 0.44 (approx) 3/7=0.43 (approx)

OpenStudy (anonymous):

^^ this is easier

hero (hero):

Yeah, if you use a calculator.

OpenStudy (anonymous):

Yes

hero (hero):

But take away the calculator...then what?

OpenStudy (anonymous):

Any ways u can divide it manually its freakin simple

hero (hero):

Of course, but I prefer cross multiplication technique. Not a fan of division and decimals.

hero (hero):

I prefer to stick with whole numbers as much as possible.

hero (hero):

The thing is, I can do the cross multiplication in my head without pencil and paper.

OpenStudy (anonymous):